# Extra dimensional axion patterns

### v1

@@ VII The only possible pattern
which illustrates how a discovery of a multiple KK signal could translate into an identification of the zero-mode in the canonical QCD axion band. Additional implications of the KK photon coupling have been discussed, for instance, in Refs. [8, 44, 45, 46]. It is also worth pointing out that, while in these regimes the massive KK modes would be decoupled from the solution to the strong CP problem, they could contribute sizably to the dark matter abundance of the Universe. This is an important modification with respect to the single axion scenario, where one focus on the parameter space where this single field makes up for the totality of the dark matter we observe. Novel analyses would be however necessary to study the thermal evolution of the plateau modes and its implications to the experimental data.
@@ Acknowledgements.
The authors would like to thank Belen Gavela for the insightful discussions that helped shape the direction of this work. We are also grateful to Javi Lizana and Pablo Quilez for the enlightening discussions, and to Enrique Fernandez, Pierluca Carenza, Jakob Moritz, and Javi Serra for their valuable input on specific aspects of this study. We also thank Peter Cox for the interesting comments on the manuscript. The authors would also like to thank Joerg Jaeckel and the Institute for Theoretical Physics of the University Heidelberg for the warm hospitality during part of the realization of this work. The work of AdG was supported by the European Union’s Horizon 2020 Marie Skłodowska-Curie grant agreement No 860881-HIDDeN and the STFC under Grant No. ST/X003167/1. This article/publication is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).

### v2

@@ VII The only possible pattern
which illustrates how a discovery of a multiple KK signal could translate into an identification of the zero-mode in the canonical QCD axion band. Additional implications of the KK photon coupling have been discussed, for instance, in Refs. [8, 46, 47, 48]. It is also worth pointing out that, while in these regimes the massive KK modes would be decoupled from the solution to the strong CP problem, they could contribute sizably to the dark matter abundance of the Universe, affecting significantly the predictions with respect to the single axion case [18, 19, 20, 27].
@@ Acknowledgements.
The authors would like to thank Belen Gavela for the insightful discussions that helped shape the direction of this work. We are also grateful to Javi Lizana and Pablo Quilez for the enlightening discussions, and to Enrique Fernandez, Pierluca Carenza, Jakob Moritz, and Javi Serra for their valuable input on specific aspects of this study. We also thank Peter Cox for the interesting comments on the manuscript and our referee, for pointing out relevant literature that was overlooked in the first version of this work. The authors would also like to thank Joerg Jaeckel and the Institute for Theoretical Physics of the University Heidelberg for the warm hospitality during part of the realization of this work. The work of AdG was supported by the European Union’s Horizon 2020 Marie Skłodowska-Curie grant agreement No 860881-HIDDeN and the STFC under Grant No. ST/X003167/1. This article/publication is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).